A father is three times as old as his son. Four years ago, he was four times as old as his son was at that time. How old is his son?
...i know it's easy but i already forget it because of chemistry my favorite subject.haha

Answer by mananth(16946)

son's age x
father's age 3x
4 years ago
son's age x-4
father's age3x-4
father 4 times older
3x-4=4(x-4)
3x-4=4x -16
-x=-16-4
-x=-12
x=12 son's age